OpenStudy (anonymous):

7(a+1)-3a=5+4(2a-1). Solve for a. I got -6 but my friend got 3/2...

OpenStudy (anonymous):

7a+7-3a=5+8a-4 4a+7=1+8a 4a+8=8a 8=4a 2=a

OpenStudy (anonymous):

did y'all distribute?

OpenStudy (anonymous):

Yes.. I did 7a+7-3a=5+... oh see i accidentally multiplied wrong, i got 6a instead of 8a.

OpenStudy (anonymous):

ah, okey doke. I love equations. They make so much sense.